# Magic dynamics in many-body localized systems

### v1

@@ Acknowledgments.
E.T. and P.S. acknowledge collaboration with X. Turkeshi and M. Dalmonte on related subjects. P.S. acknowledges insightful discussions with P. Stornati, S. Masot-Llima, A. Garcia-Saez. The work of P.R.N.F. was funded by the National Science Centre, Poland, project 2021/03/Y/ST2/00186 within the QuantERA II Programme that has received funding from the European Union Horizon 2020 research and innovation programme under Grant agreement No 101017733. P.S. acknowledges fellowship within the “Generación D” initiative, Red.es, Ministerio para la Transformación Digital y de la Función Pública, for talent atraction (C005/24-ED CV1), funded by the European Union NextGenerationEU funds, through PRTR. The work of J.Z. was funded by the National Science Centre, Poland, project 2021/43/I/ST3/01142 – OPUS call within the WEAVE programme. E. T. acknowledges support from ERC under grant agreement n.101053159 (RAVE), and CINECA (Consorzio Interuniversitario per il Calcolo Automatico) award, under the ISCRA initiative and Leonardo early access program, for the availability of high-performance computing resources and support. The study was also partially funded by the “Research Support Module” as part of the “Excellence Initiative – Research University” program at the Jagiellonian University in Kraków. We gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Centers: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2024/017289
@@ I Nonstabilizerness of Anderson states
In this section, we demonstrate how to obtain an analytical expression for the stabilizer Rényi entropy (SRE) for a perfect Anderson insulator. In this case, the $\ell$-bits do not interact with each other and, therefore, only the first term of ^$\hat{\mathcal{H}}_{\mathrm{\ell-bit}}$ remains relevant. The SRE can be rewritten as
@@
where $\mathbb{E}_{\{h\}}$ denotes the average over different disorder realizations and $\rho$ is the density matrix of the state $|\Psi(t)\rangle$. Since the system is non-interacting, the density matrix can be written as $\rho=\bigotimes_{k=1}^{L}\rho_{k}$, where $\rho_{k}$ is the reduced density matrix of the $k$-th spin. For any initial state $|\Psi\rangle=\bigotimes_{k=1}^{L}(\cos(\theta_{k}/2)|\uparrow\rangle+e^{i\phi_{k}}\sin(\theta_{k}/2)|\downarrow\rangle)$, the reduced density matrix $\rho_{k}$ at time $t$ is given by [97]:
@@
implying that to understand how the SRE grows in the non-interacting picture, it suffices to understand the dynamics of single-spin observables. The single qubit SRE can be easily obtained through the reduced density matrix $\rho_{k}$, yielding
@@
where $\delta_{k}=4h_{k}t+2\phi_{k}$. Substituting this expression into Eq. S.3, and assuming that the on-site fields $h_{k}$ are drawn independently and identically from a uniform distribution in the interval $[-W,W]$, we obtain
[…]
where the first term inside the bracket comes from the identity contribution, the second to the expectation value of $\tau^{z}$ operators and the third term is obtained from the $\tau^{x/y}$ operators. At $\theta=\pi/2$, the second term vanishes and the single-qubit SRE significantly differ from its initial value. To quantify this difference, we define the nonstabilizerness gain
Figure S.1: (a) Nonstabilizerness gain (see text) for a single qubit SRE throughout the whole parameter space ($\theta,\phi$). The greatest gain occur for an initial |+ket|$\rangle$| + ⟩ state, while the smallest occur for an initial $|T\rangle$ state. Comparison between the analytical solution given by Eq. S.5 (dashed line) and the numerical calculations of the $\ell$-bit model (solid line) for an initial (b) $|\Psi_{X}^{+}\rangle$ and (c) $|\Psi_{T}\rangle$ state. The results were obtained for $L=16$ and $10^{3}$ disorder realizations.
@@
In particular, the $X$ polarized state gives us the greatest gain. To confirm our analytical result, we consider $|\Psi_{X}^{+}\rangle=\bigotimes_{k=1}^{L}|+\rangle$ as the initial state, where |+ket|$\rangle$| + ⟩ is an eigenstate of the $\hat{X}$ operator. In this particular case, Eq. S.5 can be written as
[…]
which, in the limit $t\rightarrow\infty$, yields $\mathcal{M}_{2}\approx L\log_{2}(8/7)$. In Fig. S.1(b), we show a comparison between the analytical solution (dashed red line) and the numerical result (solid black line). The analytical solution perfectly describes the behavior of $\mathcal{M}_{2}$, showing the accuracy of our analytical calculation. In a similar spirit, we compare the exact and numerical results for an initial state $|\Psi_{T}\rangle=\bigotimes_{k=1}^{L}|T\rangle$, where $|T\rangle=(1/\sqrt{2})(|0\rangle+e^{i\pi/4}|1\rangle)$ is the magic state. In the disorderless case, $|\Psi_{T}\rangle$ has the highest SRE of a unentangled state [28]. Here, however, the initial spin precession quickly decreases its SRE value, leading to a saturation value that significantly differs from its initial value, as shown in Fig. S.1(c).
@@ II Details on the $\ell$-bit model
In the MBL regime, when the disorder is coupled to on-site operators 111 If the disorder is not coupled to on-site operators, the standard $\ell$-bit description may fail to capture the system’s physics. In such cases, a more sophisticated RSRG-X approach is required to correctly describe the model’s properties [136] , the system is simply described by the emergence of an exponentially large set of integrals of motion, also known as localized bits ($\ell$-bits). The quasilocal structure of the $\ell$-bits encapsulates key signatures of MBL, making the Hamiltonian ^$\hat{\mathcal{H}}_{\mathrm{\ell-bit}}$ (see Eq. 3 of the main letter) a natural framework to probe the nonstabilizerness dynamics in MBL. The physical degrees of freedom are related to the $\ell$-bits by quasilocal unitary transformations $U$, satisfying $\hat{\tau}_{j}^{z}=U\hat{Z}_{j}U^{\dagger}$. Ideally, $U$ reflects the system’s localized nature, with its matrix elements decaying exponentially with the distance between $\ell$-bits. However, for practical purposes, a good approximation is to take $U=I$, where $I$ is the identity operator, significantly simplifying the analysis while retaining the essential features of MBL dynamics.
@@
where $\rho_{L/2}$ is the reduced density matrix obtained by tracing out the degrees of freedom of the first half of spins. Fig. S.2(a) shows the growth of $\mathcal{S}$ starting from the initial state $|\Psi_{X}^{+}\rangle$. In the Anderson insulating case (solid black line), where interactions are absent, correlations cannot spread, and therefore $\mathcal{S}=0$. However, in interacting cases, $\mathcal{S}$ grows logarithmically in time, a hallmark of MBL. The rate of entanglement growth increases slightly as additional interaction terms are included in ^$\hat{\mathcal{H}}_{\mathrm{\ell-bit}}$ as illustrated in Fig. S.2(a). The differences between the interaction terms are more striking in the SRE behavior, as shown in Fig. S.2(b). Assuming only two-body interactions between the $\ell$-bits does not lead to the SRE reaching the Haar value, at least for the time scales analyzed here. However, as we add corrections to the $\ell$-bit Hamiltonian, the dynamics of the SRE approximates the semi-analytical prediction described in Eq. 5.
Figure S.3: (a) Nonstabilizerness spread in the $\ell$-bit model for different system sizes, with $|\Psi_{X}^{+}\rangle$ as the initial state. The dashed lines show the analytical solution (Eq. 5 of the main text), which accurately describes the SRE growth after $t\sim 1$. (b) Time evolution of $\Delta\mathcal{M}_{2}$ for different system sizes $L$. It decays polynomially in $t$ until it eventually saturates, except for $L=18$, for which the Heisenberg time is beyond the considered here. The saturation value exhibits an exponential decay with exponent $\lambda\approx-\ln{2}$, as shown in the inset.
To evaluate the accuracy of our semi-analytical description (Eq. 5 of the main letter), we examine in Fig. S.3(a) the growth of the SRE in the $\ell$-bit model for different system sizes $L$, using $|\Psi_{X}^{+}\rangle$ as the initial state. For small chains, the analytical solution (dashed lines) deviates from the behavior obtained by the numerics. However, as the system size increases, the analytical expression accurately describes the behavior of $\mathcal{M}_{2}$ at time scales where entanglement is relevant, and it saturates close to the Haar value (dotted lines) after sufficiently long times. Fig. S.3(b) further explores this convergence by showing the time evolution of the deviation from the Haar value, defined as $\Delta\mathcal{M}_{2}$. $\Delta\mathcal{M}_{2}$ decreases polynomially with time before stabilizing at a finite asymptotic value $\Delta\mathcal{M}_{2}^{\infty}$. As illustrated in the inset of Fig. S.3(b), this residual deviation decreases exponentially with system size, following a decay rate of approximately $\lambda\approx-\ln{2}$. This exponential suppression suggests that in the thermodynamic limit $\mathcal{M}_{2}$ converges to the Haar value, confirming that nonstabilizerness reaches its maximal extent in large-scale systems.
@@
III Dependence on the initial state and weight of $Z$ gates
Figure S.5: $\mathcal{W}_{Z}$ dynamics for two different states, $|\Psi_{Y}\rangle$ and $|\Psi_{Z}\rangle$, in the (a) ergodic and (b) localized regimes. In the simulation, we consider $L=16$ and average the results over, at least, 700 realizations. In the ergodic regime, the two different states evolve similarly due to the lack of integrals of motion. In the localized regime, however, the weight of $Z$ gates ($\mathcal{W}_{Z}$) depend on the initial state, leading to different saturation values of the SRE.
In the previous section, we checked the choice of initial state strongly influences the saturation value of the SRE. This behavior arises from the system’s integrals of motion: for an initial $|\Psi_{X}^{+}\rangle$ state, the expectation values of any Pauli operator containing $Z$ gates vanish, allowing the SRE to reach its maximum value. On the other hand, for a slightly perturbed eigenstate, the presence of $Z$ gates contributes nontrivially to the SRE calculation, leading to a lower saturation value. To assess how well the transverse-field Ising model (TFIM) aligns with this expected behavior, we compute the contribution of $\hat{Z}$ gates, quantified by the weight function
$\mathcal{W}_{Z}(|\Psi\rangle)=\sum_{P\in\mathcal{P}_{IZ}}\frac{|\langle\Psi|P|\Psi\rangle|^{2}}{D}$ (S.11)
where $\mathcal{P}_{IZ}$ is a subgroup of the Pauli group $\mathcal{P}_{L}$ that contains only $\hat{I}$ and $\hat{Z}$ operators. In the following analysis, we study the dynamics of $\mathcal{W}_{Z}$ for two different choices of initial state, $|\Psi_{Y}\rangle$ and $|\Psi_{Z}\rangle$, across the ergodic ($W=2.0$) and localized ($W=8.0$) regimes.
We begin by analyzing the dynamics of $\mathcal{W}_{Z}$ in the ergodic regime. In this case, the system lacks an extensive set of integrals of motion, and beyond an initial transient period, the evolution of $\mathcal{W}_{Z}$ becomes independent of the initial state, as shown in Fig. S.5(a). Here, Pauli strings containing only $Z$ gates behave similarly to other Pauli strings, allowing the SRE to reach the Haar value at sufficiently long times. The behavior changes significantly in the localized regime, as depicted in Fig. S.5(b). For an initial $|\Psi_{Y}\rangle$ state, $\mathcal{W}_{Z}$ saturates to a value close to that expected when only the identity operator contributes, indicated by the orange dashed line in the inset. This result is consistent with $\ell$-bit predictions, where expectation values of Pauli operators containing $Z$ gates should vanish for this specific state. On the other hand, for an initial $|\Psi_{Z}\rangle$ state, the weight of $Z$ gates is considerable, and it reaches a value that is significantly higher than other gates. Its behavior is almost constant in time, as expected when the system is described by a set of constants of motion. However, by fitting with a power-law decay ($\mathcal{W}_{Z}=at^{-\lambda}$), one obtains a small decay in time ($\lambda=0.005\pm 0.001$) suggesting an extremely slow delocalization at this particular disorder strength.
IV Entanglement as the “internal-clock” for the nonstabilizerness dynamics
Figure S.6: Rescaled SRE ($\mathcal{M}_{2}/f(W)$) versus entanglement entropy ($\mathcal{S}$) for the initial state $|\Psi_{R}\rangle$ and (a) $L=12$, (b) $L=14$, (c) $L=16$ and (d) $L=20$. The data collapse improves with increasing $L$, without altering the rescaling parameter $f(W)$. For $L=20$, the data were obtained via Monte-Carlo sampling using $15000$ samples, and the shaded black region indicate the sampling errors.
@@ IV Entanglement as the “internal-clock” for the nonstabilizerness dynamics
For the former choice, as shown in Fig 3(a) of the main text, both models agree without any need of rescaling by a function $f(W)$. The situation differs if we set $|\Psi_{R}\rangle$ as the initial state. In this case, in order to achieve the collapse between the curves, we minimize the distance between the results obtained for the TFIM with respect to the one obtained for the ^$\hat{\mathcal{H}}_{\mathrm{\ell-bit}}$. As shown in Fig S.6, the quality of the collapse improves as we increase the system size. Most importantly, the rescaling parameter $f(W)$ is independent of $L$, as depicted in the inset of Fig 3(b) of the main text. Hence, our phenomenological description of the SRE growth based on the integrals of motion is fully consistent with the more intricate structure of microscopic Hamiltonians.

### v2

@@ Acknowledgments.
E.T. and P.S. acknowledge collaboration with X. Turkeshi and M. Dalmonte on related subjects. P.S. acknowledges insightful discussions with P. Stornati, S. Masot-Llima, and A. Garcia-Saez. The work of P.R.N.F. and J.Z. was partially funded by the National Science Centre, Poland, project 2021/03/Y/ST2/00186 within the QuantERA II Programme (DYNAMITE) that has received funding from the European Union Horizon 2020 research and innovation programme under Grant agreement No 101017733. The work of P.R.N.F. and J.Z. was also funded by the National Science Centre, Poland, project 2021/43/I/ST3/01142 – OPUS call within the WEAVE programme. P.S. acknowledges fellowship within the “Generación D” initiative, Red.es, Ministerio para la Transformación Digital y de la Función Pública, for talent attraction (C005/24-ED CV1), funded by the European Union NextGenerationEU funds, through PRTR. E. T. acknowledges support from ERC under grant agreement n.101053159 (RAVE), and CINECA (Consorzio Interuniversitario per il Calcolo Automatico) award, under the ISCRA initiative and Leonardo early access program, for the availability of high-performance computing resources and support. The study was also partially funded by the “Research Support Module” as part of the “Excellence Initiative – Research University” program at the Jagiellonian University in Kraków. We gratefully acknowledge the Polish high-performance computing infrastructure PLGrid (HPC Centers: ACK Cyfronet AGH) for providing computer facilities and support within the computational grant no. PLG/2024/017289
@@ I Nonstabilizerness of Anderson states
In this section, we demonstrate how to obtain an analytical expression for the stabilizer Rényi entropy (SRE) for a perfect Anderson insulator. In this case, the $\ell$-bits do not interact with each other and, therefore, only the first term of $\hat{\mathcal{H}}_{\mathrm{\ell-bit}}$ remains relevant. The SRE can be rewritten as
@@
where $\mathbb{E}_{\{h\}}$ denotes the average over different disorder realizations and $\rho$ is the density matrix of the state $|\Psi(t)\rangle$. Since the system is noninteracting, the density matrix can be written as $\rho=\bigotimes_{k=1}^{L}\rho_{k}$, where $\rho_{k}$ is the reduced density matrix of the $k$-th spin. For any initial state $|\Psi\rangle=\bigotimes_{k=1}^{L}(\cos(\theta_{k}/2)|\uparrow\rangle+e^{i\phi_{k}}\sin(\theta_{k}/2)|\downarrow\rangle)$, the reduced density matrix $\rho_{k}$ at time $t$ is given by [Serbyn14]:
@@
This implies that, to understand how the SRE grows in the noninteracting picture, it suffices to understand the dynamics of single-spin observables. The single qubit SRE can be easily obtained through the reduced density matrix $\rho_{k}$, yielding
@@
where $\delta_{k}=4h_{k}t+2\phi_{k}$. Substituting this expression into Eq. S.3, and assuming that the on-site fields $h_{k}$ are drawn independently and identically from a uniform distribution in the interval [W,W][-W,W], we obtain
$\mathcal{M}_{2}=-\sum_{k=1}^{L}\frac{1}{W}\int_{0}^{W}dh\,\log_{2}\biggl{[}\,1-\frac{1}{4}\sin^{2}(2\theta_{k})\\ -\frac{1}{4}\sin^{4}\theta_{k}\sin^{2}\delta_{k}\biggr{]}.$ (S.5)
where the first term inside the bracket comes from the identity contribution, the second to the expectation value of $\tau^{z}$ operators, and the third term is obtained from the $\tau^{x/y}$ operators. At $\theta=\pi/2$, the second term vanishes and the single-qubit SRE significantly differs from its initial value. To quantify this change, we define the nonstabilizerness gain:
Figure S.1: (a) Nonstabilizerness gain (see text) for a single qubit SRE throughout the whole parameter space ($\theta,\phi$). The greatest gain occur for an initial $|+\rangle$ state, while the smallest occur for an initial $|T\rangle$ state. Comparison between the analytical solution given by Eq. S.5 (dashed line) and the numerical calculations of the $\ell$-bit model (solid line) for an initial (b) $|\Psi_{X}^{+}\rangle$ and (c) $|\Psi_{T}\rangle$ state. The results were obtained for $L=16$ and $10^{3}$ disorder realizations.
@@
In particular, the $X$ polarized state yields the greatest gain. To confirm our analytical result, we consider $|\Psi_{X}^{+}\rangle=\bigotimes_{k=1}^{L}|+\rangle$ as the initial state, where $|+\rangle$ is an eigenstate of the $\hat{X}$ operator. In this particular case, Eq. S.5 can be written as
$\mathcal{M}_{2}=-\frac{L}{W}\int_{0}^{W}dh\log_{2}\bigg{[}1-\frac{1}{4}\sin^{2}(4ht)\bigg{]}$ (S.7)
which, in the limit $t\rightarrow\infty$, yields $\mathcal{M}_{2}\approx L\log_{2}(8/7)$. In Fig. S.1(b), we show a comparison between the analytical solution (dashed red line) and the numerical result (solid black line). The analytical solution perfectly describes the behavior of $\mathcal{M}_{2}$, showing the accuracy of our analytical calculation. In a similar spirit, we compare the exact and numerical results for an initial state $|\Psi_{T}\rangle=\bigotimes_{k=1}^{L}|T\rangle$, where $|T\rangle=(1/\sqrt{2})(|0\rangle+e^{i\pi/4}|1\rangle)$ is the magic state. In the disorderless case, $|\Psi_{T}\rangle$ has the highest SRE of an unentangled state [28]. Here, however, the initial spin precession quickly decreases its SRE value, leading to a saturation value that significantly differs from its initial value, as shown in Fig. S.1(c).
@@ II Details on the $\ell$-bit model
In the MBL regime, when the disorder is coupled to on-site operators 111 If the disorder is not coupled to on-site operators, the standard $\ell$-bit description may fail to capture the system’s physics. In such cases, a more sophisticated RSRG-X approach is required to correctly describe the model’s properties [Aramthotill24], the system is simply described by the emergence of an extensive large set of integrals of motion, also known as localized bits ($\ell$-bits). The quasilocal structure of the $\ell$-bits encapsulates key signatures of MBL, making the Hamiltonian $\hat{\mathcal{H}}_{\mathrm{\ell-bit}}$ (see Eq. 3 of the main letter) a natural framework to probe the nonstabilizerness dynamics in MBL. The physical degrees of freedom are related to the $\ell$-bits by quasilocal unitary transformations $U$, satisfying $\hat{\tau}_{j}^{z}=U\hat{Z}_{j}U^{\dagger}$. Ideally, $U$ reflects the system’s localized nature, with its matrix elements decaying exponentially with the distance between $\ell$-bits. However, for practical purposes, a good approximation is to take $U=I$, where $I$ is the identity operator, significantly simplifying the analysis while retaining the essential features of MBL dynamics.
@@
where $\rho_{L/2}$ is the reduced density matrix obtained by tracing out the degrees of freedom of the first half of spins. Fig. S.2(a) shows the growth of $\mathcal{S}$ starting from the initial state $|\Psi_{X}^{+}\rangle$. In the Anderson insulating case (solid black line), where interactions are absent, correlations cannot spread, and therefore $\mathcal{S}=0$. However, in interacting cases, $\mathcal{S}$ grows logarithmically in time, a hallmark of MBL. The rate of entanglement growth increases slightly as additional interaction terms are included in $\hat{\mathcal{H}}_{\mathrm{\ell-bit}}$ as illustrated in Fig. S.2(a). The differences between the interaction terms are more striking in the SRE behavior, as shown in Fig. S.2(b). Assuming only two-body interactions between the $\ell$-bits does not lead to the SRE reaching the Haar value, at least for the time scales analyzed here. However, as we add corrections to the $\ell$-bit Hamiltonian, the dynamics of the SRE approximates the phenomenological prediction described in Eq.(5) of the main letter, as illustrated by the orange dashed line.
@@
III Entanglement as the “internal-clock” for the nonstabilizerness dynamics
Figure S.4: Rescaled SRE ($\mathcal{M}_{2}/f(W)$) versus entanglement entropy ($\mathcal{S}$) for the initial state $|\Psi_{R}\rangle$ and (a) $L=12$, (b) $L=14$, (c) $L=16$ and (d) $L=20$. The data collapse improves with increasing $L$, without altering the rescaling parameter $f(W)$. For $L=20$, the data were obtained via Monte-Carlo sampling using 1500015000 samples, and the shaded black region indicate the sampling errors.
@@ III Entanglement as the “internal-clock” for the nonstabilizerness dynamics
For the former choice, as shown in Fig 3(a) of the main text, both models agree without any need of rescaling by a function $f(W)$. The situation differs if we set $|\Psi_{R}\rangle$ as the initial state. In this case, in order to achieve the collapse between the curves, we minimize the distance between the results obtained for the TFIM with respect to the one obtained for the $\hat{\mathcal{H}}_{\mathrm{\ell-bit}}$. As shown in Fig S.4, the quality of the collapse improves as we increase the system size. Most importantly, the rescaling parameter $f(W)$ is independent of $L$, as depicted in the inset of Fig 3(b) of the main text. Hence, our phenomenological description of the SRE growth based on the integrals of motion is fully consistent with the more intricate structure of microscopic Hamiltonians.